OpenStudy (anonymous):

Graph the following lines and determine if they are parellel, perpendicular, coincide, or intersecting lines. 1. Y= 3x +2 9x-3y=-6

OpenStudy (compassionate):

Hi, Y= 3x +2 9x-3y=-6 Plug in "Y" 9x - 3(3x + 2) = -6 9x - 9x - 6 = -6 0 = 0 What does this mean?

OpenStudy (anonymous):

Can someone graph it on a graph for me

OpenStudy (compassionate):

0 = 0 cannot be graphed.

OpenStudy (compassionate):

Therefore, the lines coincide.

OpenStudy (anonymous):

Then wut line type would it be

OpenStudy (compassionate):

No line.

OpenStudy (anonymous):

Graph the following lines and determine if they are parellel, perpendicular, coincide, or intersecting lines. 2. Y=-2x+3 2x-4y=8.